OpenStudy (anonymous):

find all roots and check answer: 10x^2-5=23x

OpenStudy (anonymous):

(5x+1)(2x-5) 5x+1 = 0 or 2x-5 = 0 so x = -1/5 and x = 5/2

OpenStudy (anonymous):

awesome thanks